and this other one i don't get is::



3x+7=2x-7?
how do you do this step by step so i'll be prepared for the test tomorrow?

Answer by mananth(16946)

3x+7=2x-7
bring all the terms on one side and constants on other side.
add -7 to both sides
3x+7-7=2x-7-7
3x=2x-14
add -2x to both sides
3x-2x=2x-2x-14
x=-14
..
all the best
..
m.ananth@hotmail.ca